# A measure-valued HJB perspective on Bayesian optimal adaptive control

### v1

@@ 5.2. Subsolution property of limiting problem
where the convergence follows from Theorem 4.3. Indeed, according to this result, for any subsequence of $\mu_{k}$, there exists a further subsequence, which we still index by $k$, along which $\mathcal{W}(\xi^{v;\mu_{k}}_{1},\xi^{v;\mu}_{1})\to 0$ a.s., and along which the right-hand side of (39) thus converges to zero; since any subsequence of the original sequence admits a further subsequence with this property, the convergence must hold also along the original sequence. The same argument applies when replacing $L\varphi(\cdot,v)$ by $\beta\varphi(\cdot)$ and $k(\cdot,v)$, making use of the standing Assumption 2.1.
@@
$\liminf_{\begin{subarray}{c}h\downarrow 0\\ \mu\to\bar{\mu}\end{subarray}}\frac{1}{h}\big{(}\varphi(\mu)-\mathcal{T}_{h}\varphi(\mu)\big{)}=\lim_{k\to\infty}\frac{1}{h_{k}}\big{(}\varphi(\mu_{k})-\mathcal{T}_{h_{k}}\varphi(\mu_{k})\big{)}.$
@@
$\lim_{k\to\infty}\frac{1}{h_{k}}$ $\big{(}\varphi(\mu_{k})-\mathcal{T}_{h_{k}}\varphi(\mu_{k})\big{)}$
$\geq\limsup_{k\to\infty}\frac{1}{h_{k}}\left\{\varphi(\mu_{k})-\mathbb{E}\bigg{[}\int_{0}^{h_{k}}e^{-\beta s}k\left(\xi^{v;\mu_{k}}_{s},v\right)\mathrm{d}s+e^{-\beta h_{k}}\varphi\left(\xi^{v;\mu_{k}}_{h_{k}}\right)\bigg{]}\right\}$
$=\limsup_{k\to\infty}\frac{1}{h_{k}}\mathbb{E}\bigg{[}\int_{0}^{h_{k}}e^{-\beta s}\Big{(}\beta\varphi(\xi^{v;\mu_{k}}_{s})-k\left(\xi^{v;\mu_{k}}_{s},v\right)-L\varphi(\xi^{v;\mu_{k}}_{s},v)\Big{)}\mathrm{d}s\bigg{]}$
$\geq\liminf_{k\to\infty}\frac{1}{h_{k}}\mathbb{E}\bigg{[}\int_{0}^{h_{k}}e^{-\beta s}\Big{(}\beta\varphi(\xi^{v;\mu}_{s})-k\left(\xi^{v;\mu}_{s},v\right)-L\varphi(\xi^{v;\mu}_{s},v)\Big{)}\mathrm{d}s\bigg{]}$
$=\beta\varphi(\mu)-k\left(\mu,v\right)-L\varphi(\mu,v),$
where we used dominated convergence in the last step. Since $v\in\mathcal{U}$ was arbitrarily chosen, we may now take the supremum over $v\in\mathcal{U}$ to conclude. ∎
@@
00 0 $\geq\limsup_{n\to\infty}\frac{1}{\delta_{m_{n}}}\left\{\varphi(\mu_{n})-\inf_{v\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{\delta_{m_{n}}}e^{-\beta s}k\left(\xi^{v;\mu_{n}}_{s},v\right)\mathrm{d}s+e^{-\beta\delta_{m_{n}}}\varphi\left(\xi^{v;\mu_{n}}_{\delta_{m_{n}}}\right)\right]\right\}$
@@
$\geq H\left(\bar{\mu},\varphi(\bar{\mu}),\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(\cdot,\cdot,\bar{\mu})\right),$
@@ 5.3. Proof of main result
We first recall two results from [18] which ensure that we do indeed have comparison for the HJB equation and that the weak value function is indeed a viscosity solution of this equation. To this end, given $N$ points $x_{1},\dots,x_{N}\in\operatorname{supp}(\mu_{0})$, note that any given function $u:\mathcal{P}\to\mathbb{R}$ induces a function $\tilde{u}:\Delta^{N-1}\to\mathbb{R}$ defined by
@@
Proof.
By Lemma 4.1, the set $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$. Moreover, by the standing Assumption 2.1, $\mu\mapsto k(\mu,v)$ is continuous on $\mathcal{P}$ uniformly in $v\in\mathcal{U}$. Hence, Assumptions (i) and (ii) of [18, Theorem 9.1] are satisfied. Note first that [18, Lemma 9.2 and 9.3] still hold when relaxing their continuity assumptions on $\tilde{u},\tilde{v}$ to only imposing our (i). Moreover, by our (ii),
@@
The result can therefore be deduced along the same lines as [18, Theorem 9.1] (making obvious modifications to account for the fact that our $\mathcal{P}$ is a subset of the space considered therein). ∎
Theorem 5.11 (Theorem 6.2 in [18]).
The function $V^{weak}$ is a viscosity solution of (20).
Proof.
By Lemma 4.1, $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$, and $h(v,\cdot)$ is continuous on $\operatorname{supp}(\mu_{0})$. By the standing Assumption 2.1, we also have that $\mu\mapsto k(\mu,v)$ is continuous for every $v\in\mathcal{U}$. Hence, Assumptions (i)–(iii) of [18, Theorem 6.2] are satisfied, and the result follows. ∎
@@
It thus suffices to argue that the function $\hat{J}(u;\cdot)$ is continuous on $\mathcal{P}$ uniformly in $u\in\cup_{n\in\mathbb{N}}\mathcal{A}^{n}$; this follows however by use of the same arguments as used to prove Proposition 5.2. We next establish continuity properties of the weak value function:
@@
where we used that, by the standing Assumption 2.1, $\mu\mapsto k(\mu,v)$ is continuous, uniformly in $v\in\mathcal{U}$, and thus admits a concave and continuous modulus of continuity $w$ on the compact set $\mathcal{P}$, and the fact that $\mathcal{W}(\xi^{u;\mu},\xi^{u;\nu})$ is a submartingale (c.f. [7, Proposition 4.1.(i)]).
@@
[…]
@@
Given $\varepsilon>0$, let $n\in\mathbb{N}$ such that $V^{n}(\mu_{0})\leq V^{+}(\mu_{0})+\varepsilon$ and let $\hat{u}^{n}$ be the function provided by Corollary 5.5 for which $\hat{J}(u^{*,n};\mu_{0})\leq V^{n}(\mu_{0})+\varepsilon$ when $u^{*,n}\in\mathcal{A}^{n}$ is defined by (33) (in terms of $\hat{u}^{n}$). Define, in turn, $u^{*}$ by (21) (in terms of $\hat{u}^{n}$). Specifically, this feedback control is defined as the solution to the system (3) and (4); since the feedback function in this case is piecewise constant, it is well defined and thus belongs to $\mathcal{A}$.
We recall that $\pi^{u^{*}}_{0}=\mu_{0}$ and that $\pi^{u^{*}}$ satisfies (8). Making again use of the fact that the control is piecewise constant, according to [36, Theorem 2.1], we have pathwise uniqueness for the SDE (8) with $u$ given by (21) (viewing (8) as an equation for $\pi^{u^{*}}$ with $I^{u^{*}}$ a fixed $(\mathcal{Y}^{u^{*}},\mathbb{P})$-Brownian motion), or, equivalently, for the SDE (19) with $u$ given by (33). We therefore also have uniqueness in law (c.f. [42, proof of Theorem V.3]) and it follows that $J(u^{*})=\hat{J}(u^{*,n};\mu_{0}).$ In consequence,
@@
Next, by Theorem 5.9 and Theorem 5.11, $u=V^{+}$ and $v=V^{weak}$ are sub- and supersolutions, respectively, of (20). Moreover, by Lemma 5.12 and the preceding discussion (c.f. (41)), the continuity assumptions of Theorem 5.10 hold for $u=V^{+}$ and $v=V^{weak}$. Application of this comparison result then yields $V^{+}\leq V^{weak}$ on $\mathcal{P}$, which implies that both (31) and (42) hold as equalities. Applying Theorem 5.10 once again, making once again use of the continuity of $V^{+}$ (c.f. (41)), yields the uniqueness and continuity claims. ∎
@@ Appendix A Proof of the DPP
$\geq$’: For any $\mu\in\mathcal{P}$ and $u\in\mathcal{A}^{n}$, by use of Lemma 5.3 we obtain
@@
(43) $+e^{-\beta\delta_{n}}\int_{\Omega}\int_{0}^{\infty}e^{-\beta t}k\bigg{(}\xi^{u^{\delta_{n},\omega},\xi^{u;\mu}_{\delta_{n}}(\omega)}_{t}(\tilde{\omega}),u^{\delta_{n},\omega}_{t}(\tilde{\omega})\bigg{)}\mathrm{d}t\;\mathrm{d}\mathbb{P}(\tilde{\omega})\bigg{\}}\mathrm{d}\mathbb{P}(\omega)$
@@
$\leq$’: Let $\varepsilon>0$ and let $\{u^{\mu}\}_{\mu\in\mathcal{P}}\subset\mathcal{A}^{n}$ be a family such that
$\hat{J}(u^{\mu};\mu)\leq V^{n}(\mu)+\varepsilon/3,\quad\textrm{for every $\mu\in\mathcal{P}$.}$
From Proposition 5.2 we have uniform continuity of $V^{n}$ and $\hat{J}(u;\cdot)$, $u\in\mathcal{A}^{n}$; hence, there exists $r>0$ such that
$V^{n}(\mu)-V^{n}(\nu)\leq\varepsilon/3\quad\textrm{and}\quad\hat{J}(u^{\mu};\mu)-\hat{J}(u^{\mu};\nu)\geq-\varepsilon/3,\quad\textrm{for all $\nu\in B_{r}(\mu)$.}$
Clearly $\{B_{r}(\mu):\mu\in\mathcal{P}\}$ forms an open cover of $\mathcal{P}$ and thus there exists a finite sequence $\mu_{1},\dots\mu_{N}\in\mathcal{P}$ such that $\mathcal{P}\subset\cup_{i=1}^{N}B_{r}(\mu_{i})$. Defining
$A_{1}:=B_{r}(\mu_{1}),\qquad A_{i}:=B_{r}(\mu_{i})\setminus\{A_{1}\cup\dots\cup A_{i-1}\},\quad i=2,\dots N,$
we have that $\mathcal{P}\subset\cup_{i=1}^{N}A_{i}$ and $A_{i}\cap A_{j}=\emptyset$, $i\neq j$. Defining $u^{i}=u^{\mu_{i}}$, $i=1,\dots,N$, it then holds that
(44) $\hat{J}(u^{i};\mu)\leq\hat{J}(u^{i};\mu_{i})+\varepsilon/3\leq V^{n}(\mu_{i})+2\varepsilon/3\leq V^{n}(\mu)+\varepsilon,\quad\textrm{for all $\mu\in A_{i}$.}$
Next, let $\mu\in\mathcal{P}$ and let $u\in\mathcal{A}^{n}$ be an arbitrary control. In turn, define a control $\hat{u}\in\mathcal{A}^{n}$ by
@@
where
$\theta_{\delta_{n}}\omega(t)=\omega(\delta_{n}+t)-\omega(\delta_{n}).$
Making once again use of Lemma 5.3, and the fact that for each $i=1,\dots,N$ and $\omega\in\Omega$ such that $\xi^{u;\mu}_{\delta_{n}}(\omega)\in A_{i}$, it holds that
$\hat{u}^{\delta_{n},\omega}_{t}(\tilde{\omega})=u^{i}_{t}\left(\tilde{\omega}\right),\quad(t,\tilde{\omega})\in[0,\infty)\times\Omega,$
we obtain for $\mathbb{P}$-a.a. $\omega$,
@@
$=\int_{\Omega}\int_{0}^{\infty}e^{-\beta t}k\bigg{(}\xi^{\hat{u}^{\delta_{n},\omega},\xi^{\hat{u};\mu}_{\delta_{n}}(\omega)}_{t}(\tilde{\omega}),\hat{u}^{\delta_{n},\omega}_{t}(\tilde{\omega})\bigg{)}\mathrm{d}t\;\mathrm{d}\mathbb{P}(\tilde{\omega})$
$=\sum_{i=1}^{N}\boldsymbol{1}_{A_{i}}\left(\xi^{u;\mu}_{\delta_{n}}(\omega)\right)\left\{\int_{\Omega}\int_{0}^{\infty}e^{-\beta t}k\bigg{(}\xi^{u^{i},\xi^{u;\mu}_{\delta_{n}}(\omega)}_{t}(\tilde{\omega}),u^{i}_{t}(\tilde{\omega})\bigg{)}\mathrm{d}t\;\mathrm{d}\mathbb{P}(\tilde{\omega})\right\}$
@@
Note that, thanks to (44) and (A), for all $\mu\in\mathcal{P}$,
@@
Let $\mu\in\mathcal{P}$ and define $u\equiv u^{*,n}\in\mathcal{A}^{n}$ by (33) with respect to $\hat{u}^{n}$; it remains to argue that $u$ is ‘$\varepsilon$-optimal’. By definition, (45) holds with $\hat{u}^{n}(\mu)$ replaced by $u$. In turn, making once again use of property (45) together with the fact that $\hat{u}^{n}(\xi^{u;\mu}_{\delta_{n}}(\omega))=u^{\delta_{n},\omega}_{0},$ for each $\omega\in\Omega$, and Lemma 5.3, we obtain
@@
In consequence,
$\mathbb{E}\bigg{[}\int_{0}^{\delta_{n}}e^{-\beta t}k$ $\left(\xi^{u;\mu}_{t},u_{t}\right)\mathrm{d}t+e^{-\beta\delta_{n}}V^{n}\left(\xi^{u;\mu}_{\delta_{n}}\right)\bigg{]}$
$\geq$ $\;\mathbb{E}\bigg{[}\int_{0}^{2\delta_{n}}e^{-\beta t}k\left(\xi^{u;\mu}_{t},u_{t}\right)\mathrm{d}t+e^{-\beta 2\delta_{n}}V^{n}\left(\xi^{u;\mu}_{2\delta_{n}}\right)\bigg{]}-\varepsilon e^{-\beta\delta_{n}}.$
Repeated use of the same kind of arguments yields, for $l\in\mathbb{N}$,
$V^{n}(\mu)\geq\mathbb{E}\left[\int_{0}^{l\delta_{n}}e^{-\beta t}k\left(\xi^{u;\mu}_{t},u\right)\mathrm{d}t+e^{-\beta l\delta_{n}}V^{n}\left(\xi^{u;\mu}_{l\delta_{n}}\right)\right]-\varepsilon\big{(}1+e^{-\beta\delta_{n}}+\dots+e^{-\beta(l-1)\delta_{n}}\big{)}.$
Sending $l$ to infinity, making use of the fact that $V^{n}$ is bounded and that the geometric series converges, we may conclude. ∎

### v2

@@ 5.2. Subsolution property of limiting problem
where the convergence follows from Theorem 4.3. Indeed, according to this result, for any subsequence of $\mu_{k}$, there exists a further subsequence, which we still index by $k$, along which $\mathcal{W}(\xi^{v;\mu_{k}}_{1},\xi^{v;\mu}_{1})\to 0$ a.s., and along which the right-hand side of (41) thus converges to zero; since any subsequence of the original sequence admits a further subsequence with this property, the convergence must hold also along the original sequence. The same argument applies when replacing $L\varphi(\cdot,v)$ by $\varphi$ and $k(\cdot,v)$, making use of the standing Assumption 2.1.
@@
$\liminf_{\begin{subarray}{c}h\downarrow 0\\ \mu\to\bar{\mu}\end{subarray}}\frac{1}{h}\big(\varphi(\mu)-\mathcal{T}_{h}\varphi(\mu)\big)=\liminf_{k\to\infty}\frac{1}{h_{k}}\big(\varphi(\mu_{k})-\mathcal{T}_{h_{k}}\varphi(\mu_{k})\big).$
@@
$\liminf_{k\to\infty}$ $\frac{1}{h_{k}}\big(\varphi(\mu_{k})-\mathcal{T}_{h_{k}}\varphi(\mu_{k})\big)$
$\geq\liminf_{k\to\infty}\frac{1}{h_{k}}\left\{\varphi(\mu_{k})-\mathbb{E}\bigg[\int_{0}^{h_{k}}e^{-\beta s}k\left(\xi^{v;\mu_{k}}_{s},v\right)\mathrm{d}s+e^{-\beta h_{k}}\varphi\left(\xi^{v;\mu_{k}}_{h_{k}}\right)\bigg]\right\}$
$=\liminf_{k\to\infty}\frac{1}{h_{k}}\mathbb{E}\bigg[\int_{0}^{h_{k}}e^{-\beta s}\Big(\beta\varphi(\xi^{v;\mu_{k}}_{s})-k\left(\xi^{v;\mu_{k}}_{s},v\right)-L\varphi(\xi^{v;\mu_{k}}_{s},v)\Big)\mathrm{d}s\bigg]$
$=\liminf_{k\to\infty}\frac{1}{h_{k}}\mathbb{E}\bigg[\int_{0}^{h_{k}}e^{-\beta s}\Big(\beta\varphi(\xi^{v;\bar{\mu}}_{s})-k\left(\xi^{v;\bar{\mu}}_{s},v\right)-L\varphi(\xi^{v;\bar{\mu}}_{s},v)\Big)\mathrm{d}s\bigg]$
$=\beta\varphi(\bar{\mu})-k\left(\bar{\mu},v\right)-L\varphi(\bar{\mu},v),$
where, in order to obtain the last equality, we argued convergence by relying on the fundamental theorem of calculus for the inner integral and dominated convergence for the expected value. Since $v\in\mathcal{U}$ was arbitrarily chosen, we may now take the supremum over $v\in\mathcal{U}$ to conclude. ∎
@@
$0$ $\geq\liminf_{n\to\infty}\frac{1}{\delta_{m_{n}}}\left\{\varphi(\mu_{n})-\inf_{v\in\mathcal{U}}\mathbb{E}\left[\int_{0}^{\delta_{m_{n}}}e^{-\beta s}k\left(\xi^{v;\mu_{n}}_{s},v\right)\mathrm{d}s+e^{-\beta\delta_{m_{n}}}\varphi\left(\xi^{v;\mu_{n}}_{\delta_{m_{n}}}\right)\right]\right\}$
@@
$\geq H\left(\bar{\mu},\varphi(\bar{\mu}),\dfrac{\partial^{2}\varphi}{\partial\mu^{2}}(\cdot,\cdot;\bar{\mu})\right),$
@@ 5.3. Proof of main result
We first recall two results which ensure that the weak value function is a viscosity solution of the HJB equation and that we do have comparison for this equation. To this end, given $N$ points $x_{1},\dots,x_{N}\in\operatorname{supp}(\mu_{0})$, note that any given function $u:\mathcal{P}\to\mathbb{R}$ induces a function $\tilde{u}:\Delta^{N-1}\to\mathbb{R}$ defined by
@@
Moreover, by Lemma 4.1, $\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$ is a bounded subset of $\mathbb{R}$ and $h(v,\cdot)$ is continuous on $\operatorname{supp}(\mu_{0})$ for $v\in\mathcal{U}$. By the standing Assumption 2.1, we also have that $\mu\mapsto k(\mu,v)$ is continuous on $\mathcal{P}$ uniformly in $v\in\mathcal{U}$. Hence, Assumptions (i)–(iii) of [16, Theorem 6.2] and (i) and (ii) of [16, Theorem 9.1] are satisfied:
Theorem 5.9 (Theorem 6.2 in [16]).
The function $V^{weak}$ is a viscosity solution of (22).
@@
The formulation of this result is slightly different to that stated in [16] since we impose somewhat weaker continuity assumptions on $u,v$. In addition, we work on a compact domain. We justify these modifications as follows: Note that [16, Lemma 9.2 and 9.3] still hold when relaxing their continuity assumptions on $\tilde{u},\tilde{v}$ to only imposing our (i). Moreover, by our (ii),
@@
The theorem stated above can then be deduced along the same lines as [16, Theorem 9.1] (making obvious modifications to account for the fact that our $\mathcal{P}$ is a subset of the space considered therein).
@@
It thus suffices to argue that the function $\hat{J}(u;\cdot)$ is continuous on $\mathcal{P}$ uniformly in $u\in\cup_{n\in\mathbb{N}}\mathcal{A}^{n}$; this follows however by use of the same arguments as used to prove Proposition 5.2. In particular, $V^{+}$ satisfies the conditions on $u$ stated in Theorem 5.10. We next establish continuity properties of the weak value function:
@@
where we used that, by the standing Assumption 2.1, $\mu\mapsto k(\mu,v)$ is continuous, uniformly in $v\in\mathcal{U}$, and thus admits a concave and continuous modulus of continuity $w$ on the compact set $\mathcal{P}$, and the fact that $\mathcal{W}(\xi^{u;\mu},\xi^{u;\nu})$ is a submartingale (c.f. [8, Proposition 4.1.(i)]).
@@
where $\|\cdot\|$ denotes the Euclidean norm and $c$ is a constant depending on $T$, NN and $\sup\{h(v,x):v\in\mathcal{U},x\in\operatorname{supp}(\mu_{0})\}$. Since, with the identification between $\alpha\in\Delta^{N-1}$ and $\mu\in\mathcal{P}(\{x_{1},\dots,x_{N}\})$ given by $\mu=\alpha_{1}\delta_{x_{1}}+\dots+\alpha_{N}\delta_{x_{N}}$, the Euclidean and the Wasserstein distances are topologically equivalent, and since the obtained bound is independent of the chosen tuple, this yields continuity of $V^{weak}$ on $\mathcal{P}(\{x_{1},\dots,x_{N}\})$.
@@
Given $\varepsilon>0$, let $n\in\mathbb{N}$ such that $V^{n}(\mu_{0})\leq V^{+}(\mu_{0})+\varepsilon$ and let $\hat{u}^{n}$ be the function provided by Corollary 5.4 for which $\hat{J}(u^{*,n};\mu_{0})\leq V^{n}(\mu_{0})+\varepsilon$ when $u^{*,n}\in\mathcal{A}^{n}$ is defined by (35) in terms of $\hat{u}^{n}$. Define, in turn, the feedback control $u^{*}$ by (23) in terms of $\hat{u}^{n}$ (recall the relations (6) and (7)); since the feedback function is piecewise constant, this control is well defined and belongs to $\mathcal{A}$.
We recall that $\pi^{u^{*}}_{0}=\mu_{0}$ and that $\pi^{u^{*}}$ satisfies (13). Making again use of the fact that the control is piecewise constant, according to [33, Theorem 2.1], we have pathwise uniqueness for the SDE (13) with $u$ given by (23) (viewing (13) as an equation for $\pi^{u^{*}}$ with $I^{u^{*}}$ a fixed $(\mathcal{Y}^{u^{*}},\mathbb{P})$-Brownian motion) [^4] Or, equivalently, for the SDE (20) with $u$ given by (35).. We therefore also have uniqueness in law (c.f. [37, proof of Theorem V.3]) and it follows that $J(u^{*})=\hat{J}(u^{*,n};\mu_{0})$. In consequence,
@@
Next, by Theorem 5.8 and Theorem 5.9, $u=V^{+}$ and $v=V^{weak}$ are sub- and supersolutions, respectively, of (22). Moreover, by Lemma 5.11 and the preceding discussion (c.f. (43)), the continuity assumptions of Theorem 5.10 hold for $u=V^{+}$ and $v=V^{weak}$. Application of this comparison result then yields $V^{+}\leq V^{weak}$ on $\mathcal{P}$, which implies that both (33) and (44) hold as equalities. Applying Theorem 5.10 once again, making use of the continuity of $V^{+}$ (c.f. (43)), yields the uniqueness and continuity claims. ∎
@@ Appendix A Proof of the DPP
We first argue that a certain ‘pseudo Markov property’ holds within our context. To this end, let $n\in\mathbb{N}$ and recall the definition of $\delta_{n}$ from Section 3.2. Given a control $u\in\mathcal{A}^{n}$, we define a new control $u^{\delta_{n},\omega}\in\mathcal{A}^{n}$, for $\omega\in\Omega$, by
where
Lemma A.1.
Let $n\in\mathbb{N}$ and $u\in\mathcal{A}^{n}$. Then, for every $\mu\in\mathcal{P}$ and $\mathbb{P}$-$a.$ $\omega$,
[…]
@@
$\geq$’: For any $\mu\in\mathcal{P}$ and $u\in\mathcal{A}^{n}$, by use of Lemma A.1 we obtain
@@
(45) $+e^{-\beta\delta_{n}}\int_{\Omega}\int_{0}^{\infty}e^{-\beta t}k\bigg(\xi^{u^{\delta_{n},\omega};\xi^{u;\mu}_{\delta_{n}}(\omega)}_{t}(\tilde{\omega}),u^{\delta_{n},\omega}_{t}(\tilde{\omega})\bigg)\mathrm{d}t\;\mathrm{d}\mathbb{P}(\tilde{\omega})\bigg\}\mathrm{d}\mathbb{P}(\omega)$
@@
$\leq$’: Let $\varepsilon>0$. Thanks to the continuity properties established in Proposition 5.2, following [11], we may apply an open covering argument to obtain, for some $N\in\mathbb{N}$, sets $A_{i}\subset\mathcal{P}$, $i=1,\cdots,N$, forming a partition of $\mathcal{P}$ and controls $u^{i}\in\mathcal{A}^{n}$ such that $\hat{J}(u^{i};\mu)\leq V^{n}(\mu)+\varepsilon$, for all $\mu\in A_{i}$, $i=1,\dots N$. Now, let $\mu\in\mathcal{P}$ and let $u\in\mathcal{A}^{n}$ be an arbitrary control. In turn, define a control $\hat{u}\in\mathcal{A}^{n}$ by
@@
[…]
@@
[…]
@@
Note that, thanks to the properties of $u^{i}$, $i=1,\dots,N$, and (A), for all $\mu\in\mathcal{P}$,
@@
Let $\mu\in\mathcal{P}$ and define $u\equiv u^{*,n}\in\mathcal{A}^{n}$ by (35) with respect to $\hat{u}^{n}$; it remains to argue that $u$ is $\varepsilon$-optimal’. Making use of (46) together with the fact that $\hat{u}^{n}(\xi^{u;\mu}_{\delta_{n}}(\omega))=u^{\delta_{n},\omega}_{0},$ for each $\omega\in\Omega$, and Lemma A.1, we obtain
@@
Now, by definition, (46) holds also with $\hat{u}^{n}(\mu)$ replaced by $u$; substituting the above into this inequality, we obtain
$V^{n}(\mu)\geq$ $\;\mathbb{E}\bigg[\int_{0}^{2\delta_{n}}e^{-\beta t}k\left(\xi^{u;\mu}_{t},u_{t}\right)\mathrm{d}t+e^{-\beta 2\delta_{n}}V^{n}\left(\xi^{u;\mu}_{2\delta_{n}}\right)\bigg]-\varepsilon\big(1+e^{-\beta\delta_{n}}\big)$
$\geq$ $\;\mathbb{E}\left[\int_{0}^{l\delta_{n}}e^{-\beta t}k\left(\xi^{u;\mu}_{t},u\right)\mathrm{d}t+e^{-\beta l\delta_{n}}V^{n}\left(\xi^{u;\mu}_{l\delta_{n}}\right)\right]-\varepsilon\big(1+\dots+e^{-\beta(l-1)\delta_{n}}\big),$
where the second inequality follows by iterated use of the same arguments and holds for any $l\in\mathbb{N}$. Since $V^{n}$ is bounded and the geometric series converges, we may conclude by sending $l$ to infinity. ∎
Appendix B Technical justification of Example 2.2
In this section, we verify that the problem formulated in Example 2.2 satisfies all the conditions required in Section 2.1.
[…]
It remains to prove the following lemma which was used above:
Lemma B.1.
Let $R>1$ so that $I\subset(-R,R)$. The sensor function (9) satisfies, for some $K>0$,
(47) $\tilde{h}(\tilde{v},x)=\sum_{i=0}^{\infty}C_{i}(\tilde{v})x^{i},\quad\text{ where }\ \sum_{i=0}^{\infty}\big(R^{i}C_{i}(\tilde{v})\big)^{2}\leq K,\quad\text{for all }x,\tilde{v}\in I,$
where $)\tilde{v}\mapsto(C_{0}(\tilde{v}),C_{1}(\tilde{v}$$\ldots)$ is an injective and continuous function from $I$ to $\mathbb{R}^{\mathbb{N}}$ in the product topology.
Proof.
We can rewrite $\tilde{h}$ as
$\tilde{h}(\tilde{v},x)=e^{-\lambda x^{2}}e^{2\lambda\tilde{v}x}e^{-\lambda\tilde{v}^{2}}=e^{-\lambda\tilde{v}^{2}}\sum_{m=0}^{\infty}\frac{(2\lambda\tilde{v})^{m}}{m!}x^{m}\sum_{k=0}^{\infty}\frac{(-\lambda)^{k}}{k!}x^{2k}=\sum_{i=0}^{\infty}C_{i}(\tilde{v})x^{i},$
where
$C_{i}(\tilde{v})=e^{-\lambda\tilde{v}^{2}}\sum_{j=0}^{\lfloor\frac{i}{2}\rfloor}\frac{(-\lambda)^{j}(2\lambda\tilde{v})^{(i-2j)}}{j!(i-2j)!}.$
[…]
$e^{\lambda(x+\tilde{v})^{2}}=\sum_{i=0}^{\infty}\widetilde{C}_{i}(\tilde{v})x^{i},\quad\text{ where }\ \widetilde{C}_{i}(\tilde{v})=e^{\lambda\tilde{v}^{2}}\sum_{j=0}^{\lfloor\frac{i}{2}\rfloor}\frac{\lambda^{j}(2\lambda\tilde{v})^{(i-2j)}}{j!(i-2j)!}.$
[…]
$\sum_{i=0}^{\infty}C_{i}(\tilde{v})^{2}R^{2i}\leq\sum_{i=0}^{\infty}\frac{M^{2}}{{\widetilde{R}}^{2i}}R^{2i}=\frac{M^{2}}{1-\big(\frac{R}{\widetilde{R}}\big)^{2}}=:K,$
which completes the proof. ∎